A rectangle is three times as long as it is wide. If its length and width are both decreased by 2 cm, its are is decreased by 36 sq. cm. Find its original dimensions.
I've been having trouble with this problem every time I try to solve it I get x=-40/12. I've checked the answer but my teacher wants us to do the process right not just get the right answer.

Answer by oberobic(2304)

First, you must believe in your knowledge of the world, regardless of what the math says.
I have a formal logical proof that 1=2. That cannot be true, of course, but the proof looks good.
In this case, a rectangle cannot have negative length and width.
.
L = length in cm
W = width in cm
.
Length is three times width.
L = 3W
.
If L and W are decreased by 2 cm
L-2
W-2
.
Then its area is decreased by 36 cm^2
Original area is
A = L*W
so
(L-2)(W-2) = L*W -36
.
substitute for L=3W
.
(3W -2)(W-2) = 3W*W -36
.
3W^2 -6W -2W +4 = 3W^2 -36
.
3W^2 -8W + 4 -3W^2 = -36
.
-8W + 4 = -36
.
-8W = -40
.
W = 5 cm is width
.
L = 3W = 3(5) = 15 cm is the length
.
Check to see of the areas compare correctly.
.
L*W = 15*5 = 75 cm^2
.
(L-2)*(W-2) = 13*3 = 39 cm^2
.
75 - 39 = 36 cm^2
.
Correct.
.
Answer: The original dimensions are length = 15 cm and width = 5 cm.
.
Done.